A study showed that 67% of supermarket shoppers believe supermarket brands to be as good as national name brands. To investigate whether this result applies to its own product, the manufacturer of a national name-brand ketchup asked a sample of shoppers whether they believed that supermarket ketchup was as good as the national brand ketchup.

a. Formulate the hypotheses that could be used to determine whether the percentage of supermarket shoppers who believe that the supermarket ketchup was as good as the national brand ketchup differed from 67%.

H0: p Selectgreater than or equal to 0.67greater than 0.67less than or equal to 0.67less than 0.67equal to 0.67not equal to 0.67Item 1

Ha: p Selectgreater than or equal to 0.67greater than 0.67less than or equal to 0.67less than 0.67equal to 0.67not equal to 0.67Item 2

b. If a sample of 100 shoppers showed 51 stating that the supermarket brand was as good as the national brand, what is the p-value (to 4 decimals)?

c. At α = .05, what is your conclusion?

p-value Selectgreater than or equal to 0.05, rejectgreater than 0.05, do not rejectless than or equal to 0.05, rejectless than 0.05, rejectequal to 0.05, do not rejectnot equal to 0.05, do not rejectItem 4 H 0

d. Should the national brand ketchup manufacturer be pleased with this conclusion?
SelectYesNo
Answer #1

a)

null Hypothesis:               Ho:    p = 0.670
alternate Hypothesis:    Ha:    p 0.670

b)

51 sample success sample size 100.0 std error se =v(p*(1-p)/n) = 0.0470 sample proportion Ô x/n= 0.5100 test stat = 2 (f-p)/S\

fromn above p value =0.0006 ( please try 0.0007 if this comes wrong)

c)

p-value less than or equal to 0.05, reject

d)

Yes